# Self-Supervised Learning for Time Series Analysis: Taxonomy, Progress, and Prospects

### v1

@@ 6 Applications and Datasets
Forecasting. Time series forecasting is the process of analyzing time series data using statistics and modeling to make predictions of future windows or time points. The pretext task based on autoregressive-based forecasting is essentially a time series forecasting task. Therefore, various models based on forecasting tasks are proposed, such as Pyraformer [174], FilM [15], Quatformer [175], Informer [171], Triformer [176], Scaleformer[187], Crossformer [189], and Timesnet [172]. Moreover, we found that decomposing the series (seasonality and trend) and then learning and forecasting on the decomposed components will help improve the final forecasting accuracy, such as MICN [184] and CoST [117]. Besides, introducing an adversarial SSL is viable when missing values are in the series. For example, LGnet [200] introduces adversarial training to enhance the modeling of global temporal distribution, which mitigates the impact of missing values on forecasting accuracy.
@@
Classification and Clustering. The goal of classification and clustering tasks is the same, i.e., to identify the real category to which a certain time series sample belongs. Contrastive-based SSL methods are the most suitable choice for these two tasks since the core of contrastive learning is identifying positive and negative samples. Specifically, TS-TCC [111] introduces temporal contrast and contextual contrast in order to obtain more robust representations. TS2Vec [113] and MHCCL [133] performs a hierarchical contrastive learning strategy over augmented views, which enables robust representations. Similar to anomaly detection and prediction tasks, an adversarial-based SSL strategy can also be introduced into classification and clustering tasks. DTCR [62] propose a fake-sample generation strategy to assist the encoder in obtaining more expressive representations.
In summary, generative-based SSL is more suitable for anomaly detection and forecasting tasks, while contrastive-based SSL is more suitable for classification and clustering tasks. Adversarial-based SSL can play a role in various tasks, but in most cases, it is used as an additional regularization term to ensure that the features extracted by the model are more robust and informative. The above results and work also show that a hybrid approach (mixing multiple SSL strategies) is a better choice.
@@ 7 Discussion and Future Directions
Though promising progress has been made, the development of time series SSL is still at a developing stage. In this section, we point out some critical problems in current studies and outline several research directions worthy of further investigation.
@@ 7.1 Selection and combination of data augmentation
Data augmentation is one of the effective methods to generate augmented views in SSCL [201, 46]. The widely used methods for time series data include jitter, scaling, rotation, permutation, and warping [114, 33, 115, 116, 202]. In SimCLR [16], nine different augmentation methods for image data were discussed. The experiments show that “no single transformation suffices to learn good representations” and “the composition of random cropping and random color distortion is the most effective augmentation method”. This naturally raises the question of which one or a composition of data augmentation methods is optimal for time series SSCL. Recently, Um et al. [203] show that the combination of three basic augmentation methods (permutation, rotation, and time warping) is better than that of a single method and achieves the best performance in time series classification task. Iwana et al. [116] evaluate twelve time series data augmentation methods on 128 time series classification datasets with six different types of neural networks. The results show that the window warping and the window slicing algorithms proposed by Le et al. [204] are the most recommended algorithms. Pöppelbaum et al. [114] compared ten augmentation methods in time series classification tasks. The results show that left-to-right flipping, crop-and-resize, time warping, random smoothing, and noise appear to be the frontrunner approaches. Different evaluation frameworks give different recommendations and results. Therefore, an interesting direction is to construct a reasonable evaluation framework for time series data augmentation methods, then further select the optimal method or combination strategy.
@@ 7.3 SSL for irregular and sparse time series
Irregular and sparse time series data also widely exist in various scenarios. This data is measured at irregular time intervals, and not all the variables are available for each sample [210]. The straightforward approach to deal with irregular and sparse time series data is to use interpolation algorithms to estimate missing values [152, 153, 211]. However, interpolation-based models add undesirable noise and extra overhead to the model which usually worsens as the time series become increasingly sparse [53]. Moreover, irregular and sparse time series data is often expensive to obtain sufficient labeled data, which motivates us to build time series analysis models based on SSL in various tasks. Therefore, building SSL models directly on irregular and sparse time series data without interpolation is a valuable direction.
7.4 Theoretical analysis of contrastive learning
Contrastive learning pre-training on unlabeled data followed by supervised fine-tuning on labeled data is a widely used paradigm for learning from limited labeled data. Even the pre-training and downstream tasks have different training constraints. This has triggered many scholars to explore the theoretical basis behind SSCL. For example, Wang et al. [127] prove that the contrastive loss optimizes two key properties: alignment and uniformity, and introduce an optimizable metric to quantify each property. Wang et al. [212] build relations between the uniformity and the temperature $\tau$ in contrastive loss and point out that the temperature plays a key role in controlling the local separation and global uniformity of the embedding. Another strategy starts from the perspective of data augmentation methods and augmented views [213, 214, 215, 216]. Both Zhao et al. [214] and Wang et al. [216] point out that the transferability of contrastive learning depends mainly on the choice of data augmentation, and the learning process is to gradually cluster intra-class samples by aligning positive samples. Moreover, Tian et al. [213] start by analyzing the mutual information between augmented views and argue that the mutual information between views should be reduced while the relevance of augmented views to downstream tasks needs to be guaranteed. The existing methods provide an understanding and theoretical analysis of contrastive learning from multiple perspectives. However, there are still challenges in this area, such as the theoretical understanding of free negatives contrastive learning.
@@ 7.6 Adversarial attacks and robust analysis on time series
With the widespread use of deep neural networks in time series forecasting, classification, and anomaly detection, the vulnerability and robustness of deep models under adversarial have become a significant concern [218]. In the field of time series forecasting, Liu et al. [219] study the indirect and sparse adversarial attacks on multivariate probabilistic forecasting models for time series forecasting and propose two defense mechanisms: randomized smoothing and mini-max defense. Wu et al. [220] propose an attack strategy for generating an adversarial time series by adding malicious perturbations to the original time series to deteriorate the performance of time series prediction models. For time series classification, Karim et al. [221] focus on non-differentiable models (such as 1-NN DTW) in the time series domain and propose the student-teacher framework as a general solution to perform a proxy attack on the target model, differentiable or not. Zhuo et al. [222] summarize and compare various recent and typical adversarial attack and defense methods for fault classifiers in data-driven fault detection and classification systems, including white-box attack (FGSM[223], IGSM[224], C&\&&W attack[225], DeepFool[226]) and gray-box and black-box attack (UAP [227], SPSA [228], Random noise). The research on adversarial attacks and defenses against time series data is a worthwhile direction, but there is much less literature on this topic. Existing studies mainly involve forecasting and classification tasks. However, the impact of adversarial examples on time series self-supervised pre-training tasks is still unknown.
@@ 7.7 Time series SSL in collaborative systems
Distributed systems have been widely deployed in many scenarios, including intelligent control systems, wireless sensor networks, network file systems, the Internet of Things, etc. On the one hand, an appropriate collaborative learning strategy is fundamental in these systems, as users can train their own local models without sharing their private local data and circumvent the relevant privacy policy [229]. On the other hand, time series data is also widely distributed in various places in the system, and obtaining sufficient labeled data is also difficult, so time series SSL has great deployment potential. In recent years, federated learning has been the most popular collaborative learning framework and has been used successfully in various applications. Combining time series self-supervised learning and federated learning is a valuable research direction that can provide additional modeling tools for modern distributed systems.
7.8 Benchmark evaluation for time series SSL
SSL has many applications in time series classification, forecasting, clustering, and anomaly detection. However, most current research seeks to achieve the best performance on specific tasks and needs more discussion and evaluation of the self-supervised technique. Another interesting direction is to pay more attention to SSL, analyze its properties in time series modeling tasks, and give reliable benchmark evaluation.

### v3

@@ 6.2 Forecasting
  • Related methods. The pretext task based on autoregressive-based forecasting is essentially a time series forecasting task. Therefore, various models based on forecasting tasks are proposed, such as Pyraformer [185], FilM [15], Quatformer [186], Informer [173], Triformer [187], Scaleformer[188], Crossformer [189], and Timesnet [190]. Moreover, we found that decomposing the series (seasonality and trend) and then learning and forecasting on the decomposed components will help improve the final forecasting accuracy, such as MICN [191] and CoST [123]. Besides, introducing an adversarial SSL is viable when missing values are in the series. For example, LGnet [192] introduces adversarial training to enhance the modeling of global temporal distribution, which mitigates the impact of missing values on forecasting accuracy.
6.3 Classification and clustering
  • Task description. The goal of classification and clustering tasks is similar, i.e., to identify the real category to which a certain time series sample belongs.
@@ 6.3 Classification and clustering
  • Related methods. Contrastive-based SSL methods are the most suitable choice for these two tasks since the core of contrastive learning is identifying positive and negative samples. Specifically, TS-TCC [116] introduces temporal contrast and contextual contrast in order to obtain more robust representations. TS2Vec [118] and MHCCL [139] perform a hierarchical contrastive learning strategy over augmented views, which enables robust representations. Similar to anomaly detection and prediction tasks, an adversarial-based SSL strategy can also be introduced into classification and clustering tasks. DTCR [65] propose a fake-sample generation strategy to assist the encoder in obtaining more expressive representations.
@@ 7 Discussion and Future Directions
In this section, we point out some critical problems in current studies and outline several research directions worthy of further investigation.
@@ 7.1 Selection and combination of data augmentation
Data augmentation is one of the effective methods to generate augmented views in SSCL [193, 47]. The widely used methods for time series data include jitter, scaling, rotation, permutation, and warping [119, 45, 120, 121, 194]. In SimCLR [16], nine different augmentation methods for image data were discussed. The experiments show that “no single transformation suffices to learn good representations” and “the composition of random cropping and random color distortion is the most effective augmentation method”. This naturally raises the question of which one or composition of data augmentation methods is optimal for time series. Recently, Um et al. [195] show that the combination of three basic augmentation methods (permutation, rotation, and time warping) is better than that of a single method and achieves the best performance in time series classification task. Iwana et al. [121] evaluate twelve time series data augmentation methods on 128 time series classification datasets with six different types of neural networks. Different evaluation frameworks give different recommendations and results. Therefore, an interesting direction is to construct a reasonable evaluation framework for time series data augmentation methods, then further select the optimal method or combination strategy.
@@ 7.3 SSL for irregular and sparse time series
Irregular and sparse time series also widely exist in various scenarios. This data is measured at irregular time intervals, and not all the variables are available for each sample [201]. The straightforward approach to deal with irregular and sparse time series data is to use interpolation algorithms to estimate missing values [161, 162, 202]. However, interpolation-based models add undesirable noise and extra overhead to the model which usually worsens as the time series become increasingly sparse [54]. Moreover, irregular and sparse time series data is often expensive to obtain sufficient labeled data, which motivates us to build time series analysis models based on SSL in various tasks. Therefore, building SSL models directly on irregular and sparse time series data without interpolation is a valuable direction.
@@ 7.5 Adversarial attacks and robust analysis on time series
With the widespread use of deep neural networks in time series forecasting, classification, and anomaly detection, the vulnerability and robustness of deep models under adversarial have become a significant concern [204, 205, 206, 207]. In the field of time series forecasting, Liu et al. [205] study the indirect and sparse adversarial attacks on multivariate probabilistic forecasting models for time series forecasting and propose two defense mechanisms: randomized smoothing and mini-max defense. Wu et al. [206] propose an attack strategy for generating an adversarial time series by adding malicious perturbations to the original time series to deteriorate the performance of time series prediction models. Zhuo et al. [208] summarize and compare various recent and typical adversarial attack and defense methods for fault classifiers in data-driven fault detection and classification systems, including white-box attack (FGSM[209], IGSM[210], C&\&&W attack[211], DeepFool[212]) and gray-box and black-box attack (UAP [213], SPSA [214], Random noise). The research on adversarial attacks and defenses against time series data is a worthwhile direction, but there is much less literature on this topic. Existing studies mainly involve forecasting and classification tasks. However, the impact of adversarial examples on time series self-supervised pre-training tasks is still unknown.
7.6 Benchmark evaluation for time series SSL
SSL has many applications in time series classification, forecasting, clustering, and anomaly detection. However, most current research seeks to achieve the best performance on specific tasks and needs more discussion and evaluation of the self-supervised technique. One interesting direction is to pay more attention to SSL, analyze its properties in time series modeling tasks, and give reliable benchmark evaluation.
@@ 7.7 Time series SSL in collaborative systems
Distributed systems have been widely deployed in many scenarios, including intelligent control systems, wireless sensor networks, network file systems, etc. On the one hand, an appropriate collaborative learning strategy is fundamental in these systems, as users can train their own local models without sharing their private local data and circumventing the relevant privacy policy [215]. On the other hand, time series data is also widely distributed in various places in the system, and obtaining sufficient labeled data is also difficult, so time series SSL has great deployment potential. In recent years, federated learning has been the most popular collaborative learning framework and has been used successfully in various applications. Combining time series self-supervised learning and federated learning is a valuable research direction that can provide additional modeling tools for modern distributed systems.